OpenStudy (anonymous):

identify all equivalent forms of : sec^2(x)+3tant-11

OpenStudy (anonymous):

sec^2(x)+3tan(x)-11 sorry

OpenStudy (anonymous):

wow!!! it's too large!! how ?? replace sec^2 = tan^2+1 is 1st replace sec= 1/cos is 2nd replace tan = sin/cos but keep sec there is 3rd and so on....

OpenStudy (anonymous):

I can manipulate 11 = 1+ 10 and then 1 = sin^2 +cos^2 ha!! ridiculous question!!